OpenStudy (anonymous):

There are six chairs in a row. Person A picks one of the chairs at random. Then Person B picks a chair at random from the five chairs that remain. Find the chance that they pick chairs that are next to each other.

terenzreignz (terenzreignz):

tricky.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

conditioning on whether the chair A picks is on the leftmost or rightmost or not

OpenStudy (anonymous):

all modes of picking chairs is : \[\left(\begin{matrix}6 \\ 1\end{matrix}\right)*\left(\begin{matrix}5 \\ 1\end{matrix}\right)\]=30 and all modes of picking chairs are belong is 2*5=10 so probability I think is 1/3

OpenStudy (anonymous):

Yes, that's correct :)

OpenStudy (anonymous):

In case you need more than 1 answer: (2/6)*(1/5)+(4/6)*(2/5)=1/3 :)